OpenStudy (anonymous):

Points E(1, 2) and F(9, 8) on a coordinate grid represent side EF of triangle EFG. What is the length of side EF of the triangle? Answer 1 unit 2 units 7 units 10 units

OpenStudy (anonymous):

10

OpenStudy (anonymous):

from 1 to 9 is 8 units and from 2 to 8 is 6, so by pythagoras the distance between the points is \[\sqrt{8^2+6^2}=\sqrt{64+36}=\sqrt{100}=10\]